Dear Tutor,
I'm trying to simplify expressions using prime factors. I don't understand what that means, much less how to solve any problems. Here's an example of one of the questions:
7/30 - 29/75
(those are fractions)
If you can, please help. My math teacher doesn't explain things thoroughly. :/
Thank you so much! This is really appreciated!

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

7/30-29/75
FACTORS OF 30=2*3*5
FACTORS OF 75=3*5*5
NOW WE HAVE 2 COMMON FACTORS 3*5=15 & 2 ADDITIONAL UNIQUE FACTORS 2*5=10 WHICH WHEN MULTIPLIED TOGETHER 15*10=150 WHICH IS THE LOWEST COMMON DENOMINATOR.
(7*5-29*2)/150
(35-58)/150
-23/150 ANSWER.

Answer by stanbon(48545)

7/30 - 29/75
The least common denominator of 30 and 75 is 150
Rewrite each fraction with 150 as the denominator of each fraction:
= 7*5/150 - 29*2/150
Since the two denominators are now the same subtract to get:
= (35-58)/150
= -3/150
= -1/50
===============
Cheers,
Stan H.